# Emergent spectral geometry in the Coherence–Curvature Model

### v1

@@ I Introduction
Discrete approaches to quantum geometry and spacetime seek to understand how continuum-like structures can emerge from fundamentally combinatorial degrees of freedom. Examples include causal dynamical triangulations, group field theories, spin foam models, random geometric graphs, and a variety of network-based formulations of emergent geometry and quantum gravity [1, 2]. A central question common to these frameworks is whether ensembles of discrete configurations can exhibit effective large-scale properties that resemble those of smooth manifolds, and if so, how these properties depend on microscopic parameters and dynamics.
Two key geometric observables in this context are the spectral dimension $d_{s}$ and the Hausdorff dimension $d_{h}$. The spectral dimension is defined via diffusion processes on the underlying graph and probes how a random walker explores the configuration as a function of diffusion time. The Hausdorff dimension, by contrast, is obtained from how volumes of metric balls scale with their radius. Together, $d_{s}$ and $d_{h}$ provide complementary information about spectral and volumetric aspects of the emergent geometry, and discrepancies between them can signal nontrivial scale dependence, fractality, or multifractal behavior [1, 3, 4].
Curvature is another fundamental geometric quantity, and several notions of discrete curvature have been proposed for graphs and networks. Among these, Ollivier–Ricci curvature, defined in terms of optimal transport between local probability measures, has attracted significant attention as a flexible and computationally accessible notion of curvature on general metric spaces and graphs [5]. Ollivier–Ricci curvature has been used to analyze complex networks, study robustness and community structure [6], and explore geometric features of discrete models, with efficient implementations based on entropic-regularized optimal transport and Sinkhorn iterations [7, 8].
@@ II.2 Ollivier–Ricci curvature
We briefly recall the definition of Ollivier–Ricci curvature for graphs and the specific implementation used in the simulations [5, 6].
@@
where $W_{1}$ is the 1-Wasserstein (Earth mover’s) distance between probability measures on $V$ with respect to the graph distance $d$ [5]. On unweighted graphs, $d(u,v)=1$ for neighbors, so the denominator simply normalizes the transport distance. The vertex-wise curvature $R_{O}(v)$ appearing in Eq. (4) is taken as the average of the edge curvatures incident on $v$,
@@
The optimal transport problem defining $W_{1}(\mu_{u},\mu_{v})$ is solved numerically using an entropic-regularized scheme based on the Sinkhorn algorithm [7], as implemented in the GraphRicciCurvature Python library [6], which in turn relies on the POT (Python Optimal Transport) toolbox [8]. The entropic regularization provides a good compromise between numerical stability and computational cost for the graph sizes considered here.
@@ IV Discussion
First, at fixed couplings $(\alpha,\beta,\gamma)=(1.0,0.0091,0.15)$, we observe a robust hierarchy between the spectral dimension $d_{s}$ and the Hausdorff dimension $d_{h}$. The Hausdorff dimension $d_{h}$ converges rapidly and appears to stabilize around a value close to three. The spectral dimension $d_{s}$ increases with system size $N$, reaching values that, despite large uncertainties and evidence of incomplete convergence (Sec. III.2), consistently exceed $d_{h}$. The ratio $d_{s}/d_{h}=1.282\pm 0.399$ at $N=1024$ quantifies this hierarchy. This pattern is reminiscent of scenarios in discrete quantum gravity and random-graph models where the spectral dimension exceeds the Hausdorff dimension, signaling the presence of nontrivial connectivity and effective shortcuts that enhance diffusion [1, 3, 4].
@@
From the perspective of emergent geometry and quantum gravity, these findings illustrate how the competition in the CCM Hamiltonian can generate rich geometric behavior. The presence of a robust $d_{h}\approx 3$ combined with a tunable $d_{s}>d_{h}$ suggests that CCM-like models may realize universes with scale-dependent dimensionality, qualitatively similar to phenomena observed in other discrete approaches [1, 2].
@@ V Conclusion
Future work will aim at extending the simulations to larger volumes and deeper annealing schedules, exploring alternative curvature prescriptions and idleness parameters, and systematically mapping the $(\gamma,\beta)$ phase diagram. It would also be interesting to investigate scale-dependent effective dimensions (e.g., via diffusion-time dependent $d_{s}(t)$), to couple matter or additional degrees of freedom to the CCM graphs, and to compare the resulting geometries more directly with those arising in causal dynamical triangulations, group field theories, and other discrete approaches to quantum gravity [1, 2].

### v2

@@ I Introduction
Discrete approaches to quantum geometry and spacetime seek to understand how continuum-like structures can emerge from fundamentally combinatorial degrees of freedom. Examples include causal dynamical triangulations, group field theories, spin foam models, random geometric graphs, and a variety of network-based formulations of emergent geometry and quantum gravity [Ambjorn:2012CDT, Oriti:2011GFT]. A central question common to these frameworks is whether ensembles of discrete configurations can exhibit effective large-scale properties that resemble those of smooth manifolds, and if so, how these properties depend on microscopic parameters and dynamics.
Two key geometric observables in this context are the spectral dimension $d_{s}$ and the Hausdorff dimension $d_{h}$. The spectral dimension is defined via diffusion processes on the underlying graph and probes how a random walker explores the configuration as a function of diffusion time. The Hausdorff dimension, by contrast, is obtained from how volumes of metric balls scale with their radius. Together, $d_{s}$ and $d_{h}$ provide complementary information about spectral and volumetric aspects of the emergent geometry, and discrepancies between them can signal nontrivial scale dependence, fractality, or multifractal behavior [Ambjorn:2012CDT, Bollobas:2001RandomGraphs, Barabasi:2016book].
Curvature is another fundamental geometric quantity, and several notions of discrete curvature have been proposed for graphs and networks. Among these, Ollivier–Ricci curvature, defined in terms of optimal transport between local probability measures, has attracted significant attention as a flexible and computationally accessible notion of curvature on general metric spaces and graphs [Ollivier:2009Ricci]. Ollivier–Ricci curvature has been used to analyze complex networks, study robustness and community structure [Ni:2019Ricci], and explore geometric features of discrete models, with efficient implementations based on entropic-regularized optimal transport and Sinkhorn iterations [Cuturi:2013Sinkhorn, Flamary:2021POT].
@@ II.2 Ollivier–Ricci curvature
We briefly recall the definition of Ollivier–Ricci curvature for graphs and the specific implementation used in the simulations [Ollivier:2009Ricci, Ni:2019Ricci].
@@
where $W_{1}$ is the 1-Wasserstein (Earth mover’s) distance between probability measures on $V$ with respect to the graph distance $d$ [Ollivier:2009Ricci]. On unweighted graphs, $d(u,v)=1$ for neighbors, so the denominator simply normalizes the transport distance. The vertex-wise curvature $R_{O}(v)$ appearing in Eq. (4) is taken as the average of the edge curvatures incident on $v$,
@@
The optimal transport problem defining $W_{1}(\mu_{u},\mu_{v})$ is solved numerically using an entropic-regularized scheme based on the Sinkhorn algorithm [Cuturi:2013Sinkhorn], as implemented in the GraphRicciCurvature Python library [Ni:2019Ricci], which in turn relies on the POT (Python Optimal Transport) toolbox [Flamary:2021POT]. The entropic regularization provides a good compromise between numerical stability and computational cost for the graph sizes considered here.
@@ IV Discussion
First, at fixed couplings $(\alpha,\beta,\gamma)=(1.0,0.0091,0.15)$, we observe a robust hierarchy between the spectral dimension $d_{s}$ and the Hausdorff dimension $d_{h}$. The Hausdorff dimension $d_{h}$ converges rapidly and appears to stabilize around a value close to three. The spectral dimension $d_{s}$ increases with system size $N$, reaching values that, despite large uncertainties and evidence of incomplete convergence (Sec. III.2), consistently exceed $d_{h}$. The ratio $d_{s}/d_{h}=1.282\pm 0.399$ at $N=1024$ quantifies this hierarchy. This pattern is reminiscent of scenarios in discrete quantum gravity and random-graph models where the spectral dimension exceeds the Hausdorff dimension, signaling the presence of nontrivial connectivity and effective shortcuts that enhance diffusion [Ambjorn:2012CDT, Bollobas:2001RandomGraphs, Barabasi:2016book].
@@
From the perspective of emergent geometry and quantum gravity, these findings illustrate how the competition in the CCM Hamiltonian can generate rich geometric behavior. The presence of a robust $d_{h}\approx 3$ combined with a tunable $d_{s}>d_{h}$ suggests that CCM-like models may realize universes with scale-dependent dimensionality, qualitatively similar to phenomena observed in other discrete approaches [Ambjorn:2012CDT, Oriti:2011GFT].
@@ V Conclusion
Future work will aim at extending the simulations to larger volumes and deeper annealing schedules, exploring alternative curvature prescriptions and idleness parameters, and systematically mapping the $(\gamma,\beta)$ phase diagram. It would also be interesting to investigate scale-dependent effective dimensions (e.g., via diffusion-time dependent $d_{s}(t)$), to couple matter or additional degrees of freedom to the CCM graphs, and to compare the resulting geometries more directly with those arising in causal dynamical triangulations, group field theories, and other discrete approaches to quantum gravity [Ambjorn:2012CDT, Oriti:2011GFT].